OpenStudy (soahelp):

medal Jason left a bin outside in his garden to collect rainwater. He notices that 1 over 5 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin.

OpenStudy (soahelp):

@igreen

OpenStudy (texaschic101):

(1/5) / (2/3) = x/1 --- (1/5 water to 2/3 bin = x water to 1 bin) now we cross multiply 2/3x = 1/5 x = (1/5) / (2/3) x = 1/5 * 3/2 x = 3/10 check.. 1/5 = 2/3 of 3/10 1/5 = 2/3 * 3/10 1/5 = 6/30 1/5 = 1/5 (correct) if I am not mistaken, 3/10 gallons of water fills the entire bin.....but this is a proportion...not exactly an equation.

OpenStudy (texaschic101):

I guess the equation could be 2/3x = 1/5 with x representing 1 entire bin